I have a tough time with word problems.
A rectangular parking lot is 50ft longer than it is wide. Determine the dimensions of the parking lot if it measures 250 ft diagonally

Answer by checkley77(12844)

L=W+50
L^2+W^2=250^2
(W+50)^2+W^2=62,500
W^2+100W+2,500+W^2=62,500
2W^2+100W+2,500-62,500=0
2W^2+100W-60,000=0
2(W^2+50W-30,000)=0
2(W+200)(W-150)=0
W-150=0
W=150 ANS. FOR THE WIDTH.
L=150+50=200 ANS. FOR THE LENGTH.
PROOF:
200^2+150^2=250^2
40,000+22,500=62,500
62,500=62,500